OpenStudy (anonymous):

Which of the following is the simplified form of the quantity 4 y squared minus 1 over the quantity 2 y squared plus 3 y minus 2 ? 2y minus 1 over y minus 2 , with the restrictions y ≠ –1 over 2 and y ≠ –2 2y plus 1 over y plus 2 , with the restrictions y ≠ 1 over 2 and y ≠ 2 2y minus 1 over y minus 2, with the restriction y ≠ –1 over 2 and y ≠ 2 2y plus 1 over y plus 2, with the restriction y ≠ 1 over 2 and y ≠ –2

OpenStudy (anonymous):

\[\frac{ 4y^2-1 }{ 2y^2+3y-2 }\]

OpenStudy (luigi0210):

Doesn't look that bad >.> Do you know how factor using the difference of perfect squares?

OpenStudy (anonymous):

No?...

OpenStudy (luigi0210):

a^2-b^2=(a+b)(a-b)

OpenStudy (anonymous):

Oh god, I'll be back. My head hurts.

OpenStudy (luigi0210):

Sorry :/

OpenStudy (anonymous):

it is okay.

OpenStudy (luigi0210):

I hope you feel better!

OpenStudy (luigi0210):

Here I'll do the first part for you: \[\frac{ 4y^2-1 }{ 2y^2+3y-2 }=\frac{ (2y-1)(2y+1) }{ (2y-1)(y+2) }\]